two coherent monochromatic light beams of intensities I and 4I are superposed;the maximum and minimum possible intensities in the resulting beam are

Grade:12

1 Answers

Arun
6 years ago
Dear Deepika
 
intensity is directly proportional to square of 
amplitude 
let I=k(A)^2 
4I=k(B)^2 
where A & B are amplitudes 
k is a constant 
then A+B gives maximum possible amplitude and A-B gives minimum 
I=sqrt(k)*A 
2I=sqrt(k)*B 
add both of them and also subtract you get 
9I=k(a+b)^2 
I=k(a-b)^2 
therefore max and min are 9I and I
 
 
Regards
Arun (askIITians forum expert)